OpenStudy (anonymous):

Let y= 3 t^5 - 3. Find D, E, and F so that (d^2 y)/(d t^2) = Dt^E +F

OpenStudy (anonymous):

D=60 F=0

OpenStudy (campbell_st):

Differentiate \[f'(t) = 15t^4 \] \[f"(t) = 60t^3\] D = 60, E = 3 and F = 0

OpenStudy (anonymous):

what if its (dy)/(dt)|_(t = 2) =

OpenStudy (anonymous):

15*2^4

OpenStudy (anonymous):

so it would be 240?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

if so how could I find the 2nd deriv? from 240

OpenStudy (campbell_st):

you can't find the 2nd derivative.... in terms of t before doing anything

OpenStudy (anonymous):

so I would have to go back and plug 2 into the 2nd deriv equation?

OpenStudy (anonymous):

for t

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

ok thanx